OpenStudy (anonymous):

how do i solve 7x(2) + 30 = 9

OpenStudy (anonymous):

7x(2) is 7 x square

OpenStudy (anonymous):

is this \[7x^2+30=9\]?

OpenStudy (anonymous):

yes 7x squared

OpenStudy (anonymous):

then \[7x^2=9-30=-21\] \[x^2=-3\] and you are either stuck if you are using real numbers or else you get two complex solutions

OpenStudy (anonymous):

7x^2 -21 =0 7(x^2-3) =0 x^2 -3 =0 x^2 = 3 x = + sqrt (3) or x = - sqrt(3)

OpenStudy (anonymous):

ok i got both of your answers thanks guys

OpenStudy (anonymous):

@rita i think it should be \[7x^2=-21\] rather than \[7x^2-21=0\]

OpenStudy (anonymous):

yeah